# System Design of the Ultra Mobility Vehicle: A Driving, Balancing, and Jumping Bicycle Robot

### v1

@@ Introduction
Legged robots have demonstrated fantastic capabilities in recent years. By selecting discrete footholds, legged systems are able to handle vertical discontinuities that often pose challenges for wheeled platforms. For example, the humanoid Atlas [7] demonstrated the ability to jump onto boxes, bound between slanted surfaces, and even vault over obstacles. The quadruped ANYmal is able to climb over boxes that are taller than its standing height [27]. However, the leg morphologies that give these robots their impressive abilities are quite complex: Bipedal robots typically have six degrees of freedom (DoFs) per leg [12, 7] while quadrupeds tend to have three DoFs per leg [28, 5, 30]. Therefore, both bipeds and quadrupeds typically have twelve actuated DoFs in the legs, and as DoF count increases, so does mass, manufacturing cost, computational complexity, and risk of mechanical failure. In contrast, wheeled robots can have as few as two actuated DoFs [17, 61, 18] and excel at fast, energy-efficient transportation on relatively smooth terrain, but have difficulty overcoming obstacles. Efforts to address these tradeoffs have given rise to the “wheel-legged hybrid,” where the feet of a legged robot are replaced with driven wheels [31, 4, 63]. Such robots have undeniable benefits; for example, Bjelonic and colleagues found that, compared to walking, wheels halved the cost of transport (CoT) [4]. Going further, several works offset the cost of the additional wheel DoFs by removing leg DoFs. For instance, Ascento [34] has just four actuated DoFs. Other examples with simplified leg mechanisms include Diablo [39] and Handle [6], both of which have demonstrated the ability to hop and traverse stairs. These wheel-legged hybrids use a multi-track wheel configuration (where wheels are arranged side-by-side) as opposed to a single-track wheel configuration (where wheels are arranged in tandem, like a bicycle or scooter). While multi-track systems are easier to stabilize, they require additional suspension mechanisms to handle roll and deflection caused by uneven terrain and dynamic motion. In addition, multi-track systems have an inherently wider footprint, and thus have more difficulty with tight spaces and small affordances. As such, they have greater mass, greater mechanical complexity, and reduced maneuverability.
@@
In this paper, we present the design of the UMV system, its capabilities, and the control and state estimation methods employed on the platform to push the limits of robotic athleticism. We first present an overview of the hardware design (Figs. 1 and 2). Next, we showcase several experiments demonstrating UMV’s capabilities, including driving, turning in-place and track-standing (Movie 1), wheelies and hopping (Fig. 3), table jumps (Fig. 4), and flipping (Fig. 5). We then explain how the morphology of the robot was designed and optimized (Figs. 6, 7, and S4). Finally, we cover details of the state estimation and RL frameworks. The nomenclature used in this article is tabulated in Supplementary Table S1.
List of Movies 1 UMV combines the bicycle form with legged agility, allowing for athletic motions, including: (1) track-standing, in which the robot balances in-place using only wheel and steering actuation; (2) backward driving, (3) “shimmy-turning,” in which the robot lifts the front wheel to yaw in-place about the rear wheel, (4) lateral hopping on a single wheel, (5) front flipping, and (6) high-clearance table jumping.
@@ Results
The Head is composed of an aluminum 7075-T6 structural frame. It houses the jumping actuators, as well as compute, batteries, and most of the robot’s electronics. These are protected by a roll cage of carbon-fiber reinforced polymer (CFRP) plates and tubes, with an outer layer of foam to damp impacts. The Neck is machined from 7075-T6 in two halves. Like the Head, its “clamshell” structure maximizes cross-sectional second moment of area per unit mass, improving stiffness against lengthwise torsion and lateral forces. An additional benefit is that electrical cabling is routed through the hollow center. The Bike link consists of a magnesium AZ31B-H24 structural element clamped over the down tube of a children’s CFRP push bike (specifically, a Specialized Hotwalk Carbon). The magnesium structure houses the steering and rear wheel drive actuators as well as the communication and power distribution electronics. The Hotwalk’s CFRP fork and wheel rims are also used, where the output pulley of the rear drive belt is bonded to the rear rim. Additional struts reinforce the the rear stays to prevent deflection from loosening the rear wheel belt drive during takeoff and landing.
@@
In this experiment, we tested the robot’s ability to drive and balance dynamically. The driving behavior tracks the forward velocity and heading commands (coming from either a joystick or higher-level planner) as the robot moves across flat ground. While bicycles are known to have some passive stability while moving forward, backward movement is inherently unstable [43]. Movie 1 shows, though, that a learned policy allows UMV to maintain its balance driving both forwards and backwards. The robot is also able to turn in place about a (stationary) rear wheel contact point. We call this the “shimmy-turn”, where the robot uses the reaction mass of the Head to quickly step the front wheel laterally, creating yaw about the rear wheel. Rapidly repeating this pattern executes in-place turning. This behavior is differentiated from small radius turns with both wheels in ground contact, which requires a small amount of forward motion. The “shimmy-turn” behavior emerged from the learned policy as a viable control strategy without being explicitly designed. Finally, UMV is capable of holding a “track-stand.” This behavior is basically balancing in place, over the contact line between the front and rear wheels, with very little motion. It can be performed using just fork and rear wheel actuation, while adding Head actuation further improves balance.
@@ Discussion
The jump policies have two extra observations: a commanded goal location on the terrain in the world frame as seen from above $\in\penalty 10000\ \mathbb{R}^{2}$, and a vector of a robot-relative heights observing a local grid on terrain elevations in the direction of travel $\in\penalty 10000\ \mathbb{R}^{112}$. The flip policy has an extra observation, the phase variable $\Theta\in\mathbb{R}^{1}$, that increases linearly from 0 to 1 over the duration of a reference trajectory.
@@ Acknowledgments
The authors would like to thank Abel Gawel, Aleksandrs Ecins, Ananya Agarwal, Katie Hughes, Luca Buonocore, Marco Zenti, Petr Listov, and Victor Reijgwart for their technical contributions, dedicated support, and insightful discussions throughout the development and testing of the UMV platform. Funding: This research is supported by the Robotics and AI (RAI) Institute, which, in turn, receives funding from Hyundai Motor Group. The opinions expressed herein are those of the authors and independent of the views of the funding organizations. Author contributions: The Core Contributors developed and implemented the design optimization pipeline, designed and assembled the UMV robot, developed the embedded and training-tool software, trained most of the policies, and oversaw the hardware experiments on robots. They also contributed to the writing of the manuscript. They assisted with engineering tasks to support the project, including deployment code on hardware, data collection and curation, as well as supporting hard-ware experiments. The Project Leads set the overarching vision by defining the core scientific goals and managed the project. Competing interests: The authors attest that there are no competing interests to declare. Data and materials availability: Data to support the paper’s conclusions are present herein or the Supplementary Materials.

### v2

# arXiv:2602.22118v2[cs.RO] 27 Feb 2026

@@ Introduction
Legged robots have demonstrated fantastic capabilities in recent years. By selecting discrete footholds, legged systems are able to handle vertical discontinuities that often pose challenges for wheeled platforms. For example, the humanoid Atlas [7] demonstrated the ability to jump onto boxes, bound between slanted surfaces, and even vault over obstacles. The quadruped ANYmal is able to climb over boxes that are taller than its standing height [27]. However, the leg morphologies that give these robots their impressive abilities are quite complex: Bipedal robots typically have six degrees of freedom (DoFs) per leg [12] while quadrupeds tend to have three DoFs per leg [28, 5, 30]. Therefore, both bipeds and quadrupeds typically have twelve actuated DoFs in the legs, and as DoF count increases, so does mass, manufacturing cost, computational complexity, and risk of mechanical failure. In contrast, wheeled robots can have as few as two actuated DoFs [17, 61, 18] and excel at fast, energy-efficient transportation on relatively smooth terrain, but have difficulty overcoming obstacles. Efforts to address these tradeoffs have given rise to the “wheel-legged hybrid,” where the feet of a legged robot are replaced with driven wheels [31, 4, 63]. Such robots have undeniable benefits; for example, Bjelonic and colleagues found that, compared to walking, wheels halved the cost of transport (CoT) [4]. Going further, several works offset the cost of the additional wheel DoFs by removing leg DoFs. For instance, Ascento [34] has just four actuated DoFs. Other examples with simplified leg mechanisms include Diablo [39] and Handle [6], both of which have demonstrated the ability to hop and traverse stairs. These wheel-legged hybrids use a multi-track wheel configuration (where wheels are arranged side-by-side) as opposed to a single-track wheel configuration (where wheels are arranged in tandem, like a bicycle or scooter). While multi-track systems are easier to stabilize, they require additional suspension mechanisms to handle roll and deflection caused by uneven terrain and dynamic motion. In addition, multi-track systems have an inherently wider footprint, and thus have more difficulty with tight spaces and small affordances. As such, they have greater mass, greater mechanical complexity, and reduced maneuverability.
@@
In this paper, we present the design of the UMV system, its capabilities, and the control and state estimation methods employed on the platform to push the limits of robotic athleticism. We first present an overview of the hardware design (Figs. 1 and 2). Next, we showcase several experiments demonstrating UMV’s capabilities, including driving, turning in place and track-standing (Movie 1), wheelies and hopping (Fig. 3), table jumps (Fig. 4), and flipping (Fig. 5). We then explain how the morphology of the robot was designed and optimized (Figs. 6, 7, and S4). Finally, we cover details of the state estimation and RL frameworks. The nomenclature used in this article is tabulated in Supplementary Table S1.
List of Movies 1 UMV combines the bicycle form with legged agility, allowing for athletic motions, including: (1) track-standing, in which the robot balances in place using only wheel and steering actuation; (2) backward driving, (3) “shimmy-turning,” in which the robot lifts the front wheel to yaw in place about the rear wheel, (4) lateral hopping on a single wheel, (5) front flipping, and (6) high-clearance table jumping.
@@ Results
The Head is composed of an aluminum 7075-T6 structural frame. It houses the jumping actuators, as well as compute, batteries, and most of the robot’s electronics. These are protected by a roll cage of carbon-fiber reinforced polymer (CFRP) plates and tubes, with an outer layer of foam to damp impacts. The Neck is machined from 7075-T6 in two halves. Like the Head, its “clamshell” structure maximizes cross-sectional second moment of area per unit mass, improving stiffness against lengthwise torsion and lateral forces. An additional benefit is that electrical cabling is routed through the hollow center. The Bike link consists of a magnesium AZ31B-H24 structural element clamped over the down tube of a children’s CFRP push bike (specifically, a Specialized Hotwalk Carbon). The magnesium structure houses the steering and rear wheel drive actuators as well as the communication and power distribution electronics. The Hotwalk’s CFRP fork and wheel rims are also used, where the output pulley of the rear drive belt is bonded to the rear rim. Additional struts reinforce the rear stays to prevent deflection from loosening the rear wheel belt drive during takeoff and landing.
@@
In this experiment, we tested the robot’s ability to drive and balance dynamically. The driving behavior tracks the forward velocity and heading commands (coming from either a joystick or higher-level planner) as the robot moves across flat ground. While bicycles are known to have some passive stability while moving forward, backward movement is inherently unstable [43]. Movie 1 shows, though, that a learned policy allows UMV to maintain its balance driving both forwards and backwards. The robot is also able to turn in place about a (stationary) rear wheel contact point. We call this the “shimmy-turn”, where the robot uses the reaction mass of the Head to quickly step the front wheel laterally, creating yaw about the rear wheel. Rapidly repeating this pattern executes in place turning. This behavior is differentiated from small radius turns with both wheels in ground contact, which requires a small amount of forward motion. The “shimmy-turn” behavior emerged from the learned policy as a viable control strategy without being explicitly designed. Finally, UMV is capable of holding a “track-stand.” This behavior is basically balancing in place, over the contact line between the front and rear wheels, with very little motion. It can be performed using just fork and rear wheel actuation, while adding Head actuation further improves balance.
@@ Discussion
The jump policies have two extra observations: a commanded goal location on the terrain in the world frame as seen from above $\in\penalty 10000\ \mathbb{R}^{2}$, and a vector of robot-relative heights observing a local grid on terrain elevations in the direction of travel $\in\penalty 10000\ \mathbb{R}^{112}$. The flip policy has an extra observation, the phase variable $\Theta\in\mathbb{R}^{1}$, that increases linearly from 0 to 1 over the duration of a reference trajectory.
@@ Acknowledgments
The authors would like to thank Abel Gawel, Aleksandrs Ecins, Ananya Agarwal, Jessica Hodgins, Katie Hughes, Luca Buonocore, Marco Zenti, Petr Listov, and Victor Reijgwart for their technical contributions, dedicated support, and insightful discussions throughout the development and testing of the UMV platform.
Funding: This research is supported by the Robotics and AI (RAI) Institute, which, in turn, receives funding from Hyundai Motor Group. The opinions expressed herein are those of the authors and independent of the views of the funding organizations.
Author contributions: The Core Contributors developed and implemented the design optimization pipeline, designed and assembled the UMV robot, developed the embedded and training-tool software, trained most of the policies, and oversaw the hardware experiments on robots. They also contributed to the writing of the manuscript. They assisted with engineering tasks to support the project, including deployment code on hardware, data collection and curation, as well as supporting hardware experiments. The Project Leads set the overarching vision by defining the core scientific goals and managed the project.
Competing interests: The authors attest that there are no competing interests to declare.
Data and materials availability: Data to support the paper’s conclusions are present herein or the Supplementary Materials.